OpenStudy (anonymous):

What is the solution to the system of equations? 4x + y = 2 x - y = 3

OpenStudy (anonymous):

x=1, y=-2

OpenStudy (anonymous):

Solve the second equation for x and replace it on the first one. x=3-y 4 (3-y)+y=2, 12+4y+y=2, 5y=-10 , y=-10/5=-2. plug this into the equation and find the value of x